### v1

# Community-Based Early-Stage Chronic Kidney Disease Screening using Explainable Machine Learning for Low-Resource Settings

Abstract
Background: Early detection of chronic kidney disease (CKD) is essential for preventing progression to end-stage renal disease. However, existing screening tools—primarily developed using populations from high-income countries—often underperform in Bangladesh and South Asia, where risk profiles differ. Most of these tools rely on simple additive scoring functions and are based on data from patients with advanced-stage CKD. Consequently, they fail to capture complex interactions among risk factors and are limited in predicting early-stage CKD.
Objective: Develop and evaluate an explainable machine learning (ML) framework for community-based early-stage CKD screening for low-resource settings, tailored to the Bangladeshi and South Asian population context.
Methods: We used a community-based dataset from Bangladesh, the first such CKD dataset in South and South Asia, and evaluated twelve ML classifiers across multiple feature domains. Ten complementary feature selection techniques were applied to identify robust, generalizable predictors. The final models were assessed using 10-fold cross-validation. External validation was conducted on three independent datasets from India, the UAE, and Bangladesh. SHAP (SHapley Additive exPlanations) was used to provide model explainability.
Results: An ML model trained on an RFECV-selected feature subset achieved a balanced accuracy of 90.40%, whereas minimal non-pathology-test features demonstrated excellent predictive capability with a balanced accuracy of 89.23%, often outperforming larger or full feature sets. Compared with existing screening tools, the proposed models achieved substantially higher accuracy and sensitivity while requiring fewer and more accessible inputs. External validation confirmed strong generalizability with 78% to 98% sensitivity. SHAP interpretation identified clinically meaningful predictors consistent with established CKD risk factors.
Conclusions: Accurate, interpretable, and scalable early-stage CKD screening is achievable using only non-pathology-test features. This framework offers a practical pathway for community-level CKD screening in resource-constrained settings where laboratory testing is limited.
keywords:
Chronic kidney disease , early stage , machine learning , explainable , feature selection , community-based , screening
journal:
[…]
[3]organization=Health Systems and Population Studies Division, International Centre for Diarrhoeal Disease Research, Bangladesh (icddr,b), city=Dhaka, postcode=1212,country=Bangladesh
[…]
[4]organization=Technical Training Unit, International Centre for Diarrhoeal Disease Research, Bangladesh (icddr,b), city=Dhaka, postcode=1212,country=Bangladesh
1 Introduction
The kidneys play a vital role in maintaining physiological balance within the human body. Their primary functions include filtering waste products and excess fluids from the blood, regulating blood pressure through renin-angiotensin pathways, maintaining acid-base balance, managing electrolyte levels, and stimulating red blood cell production through the secretion of erythropoietin [lote1994principles]. Chronic kidney disease (CKD) is a progressive condition marked by the gradual decline in kidney function, eventually requiring renal replacement therapy such as dialysis or kidney transplantation. Clinically, CKD is diagnosed by persistent markers of kidney damage or a sustained reduction in the estimated glomerular filtration rate (eGFR) below 60 mL/min/1.73 $m^{2}$ for more than three months [levin2013kidney].
CKD represents a critical and growing global health challenge, with profound socioeconomic implications. Globally, kidney disease affects approximately 850 million people [jager2019single], with CKD estimated to have a prevalence of 11–13% [hill2016global]. In 2017, around 1.2 million people died from CKD and it was the 12th leading cause of death worldwide [bikbov2020global]. The overall number of CKD-related deaths has increased substantially over the recent decade, yet resources for managing this silent epidemic remain limited in most developing regions [bikbov2020global]. In low- and middle-income countries (LMICs), including South Asia, the burden is particularly severe due to limited healthcare infrastructure, shortage of nephrology care, and poor access to diagnostics [abraham2016chronic]. In Bangladesh, community-based studies in urban and rural populations suggest a prevalence rate of CKD 26% and 22%, respectively [anand2014high, sarker2021community, banik2021prevalence], with most patients unaware of their condition [iqbal2018knowledge].
Early-stage CKD detection is essential to delay or prevent the onset of end-stage kidney disease (ESKD), reduce cardiovascular complications, and avoid the substantial costs associated with renal replacement therapies (RRT). Evidence suggests that CKD progression can be slowed considerably through early lifestyle modification, pharmacological management of comorbidities, and patient education [pollock2024framework]. However, early-stage CKD is typically asymptomatic; only about 4% of individuals with CKD stages 1–2 and fewer than 10% with stage 3 are aware of their condition [plantinga2008patient], and most patients are diagnosed only after clinical symptoms emerge, by which time renal impairment is often irreversible [levey2022chronic]. The failure to detect CKD at an early stage not only worsens patient outcomes but also imposes a significant economic burden on health systems [luyckx2021global], particularly in low-income settings where RRT is prohibitively expensive and largely inaccessible [niang2018hemodialysis].
Standard clinical diagnosis of CKD relies on laboratory or pathology tests such as serum creatinine levels, estimated glomerular filtration rate (eGFR), and the urinary albumin-to-creatinine ratio (ACR) [levin2024executive]. These assessments require access to well-equipped diagnostic facilities, trained personnel, and regular follow-up—resources that are scarce in many low-resource settings [stanifer2014chronic]. Early diagnosis is particularly challenging in the Global South, including Bangladesh, where a large proportion of the population lives in rural and peri-urban communities with limited access to healthcare infrastructure, diagnostic laboratories, and nephrology specialists [kabir2023capacity]. In addition, poor awareness among patients and providers, inadequate communication between them, low risk perception among high-risk individuals, and the financial burden associated with travel, laboratory testing, and clinical consultations further deters early CKD diagnosis [zeba2020early, jafar2020access, nazar2014barriers, kahn2015chronic]. Together, these structural and socioeconomic barriers contribute to delayed recognition of CKD and missed opportunities for timely intervention.
Several community-based CKD screening tools have been developed over the past two decades to facilitate risk assessment [stolpe2022external]; however, most were designed in high-income countries and exhibit limited relevance for low-resource settings (Table 2 provides a summary). Prominent tools include the SCORED [bang2007development] from the United States , kshirsagar2008simple based on the U.S. population, the thakkinstian2011simple developed in Thailand, the kearns2013predicting from England, and kwon2019simple from South Korea. These tools rely largely on demographic and medical history variables such as age, sex, diabetes, hypertension, and cardiovascular disease, with several incorporating clinical examination features (e.g., anemia, blood pressure, dipstick proteinuria). Importantly, most were designed to detect CKD stages 3–5, thereby overlooking early-stage disease where intervention is most impactful. Their dependency on Western cohorts also limits applicability to populations in the Global South, where environmental exposures, health system constraints, and sociodemographic risk patterns differ markedly. Moreover, most models employ simple additive scoring functions, restricting their ability to capture complex interactions among risk factors. These limitations underscore the need for context-appropriate, low-cost, and explainable machine learning approaches tailored to early-stage CKD screening in resource-constrained environments.
Machine learning (ML)–based predictive tools, particularly those relying on easily obtainable features such as age, blood pressure, and medical history, offer a promising avenue for improving CKD risk assessment in low-resource settings. A substantial number of studies have applied ML techniques for CKD identification, and comprehensive summaries of these works are available in several review articles [sanmarchi2023predict, delrue2024application, gogoi2025machine]. Notably, the vast majority of ML studies are based on the UCI-2015 CKD dataset [chronic_kidney_disease_336], which we also used for external validation [gogoi2025machine]. However, many of these studies include pathology-test-based markers such as serum creatinine or eGFR as input features for model development (e.g., [gogoi2025chronic, dharmarathne2024diagnosis, islam2023chronic, ghosh2025machine, ghosh2024investigation]). This practice introduces a form of information leakage and circular reasoning, as these biomarkers are themselves the clinical criteria used to diagnose CKD; if such values are already available, an ML model becomes redundant. Moreover, pathology tests such as serum creatinine and eGFR are often inaccessible in rural or peri-urban areas and remain financially prohibitive for many individuals in the Global South. Consequently, these ML studies are unsuitable for community-based CKD screening and cannot be deployed in resource-limited populations. To the best of our knowledge, no prior ML study has focused specifically on community-based early-stage CKD screening while performing extensive feature selection, rigorous external validation across multiple datasets, and systematic comparison with established clinical screening tools as conducted in this work.
2 Methodology
Figure 1 provides an overview of our methodological pipeline. The workflow consists of several steps: (i) dataset preprocessing, including imputation, discretization, harmonization and one-hot encoding; (ii) feature ranking and selection using statistical, regularization-based, and wrapper-based methods; (iii) model development through training a diverse set of machine learning classifiers; (iv) performance evaluation, incorporating stratified cross-validation, benchmarking against state-of-the-art (SOTA) methods, and external validation on independent datasets; and (v) model explainability using SHAP to interpret both global and individual-level predictions. Together, these steps establish a rigorous and interpretable framework for early CKD prediction.
Figure 1: A schematic overview of our methodology
2.1 Datasets
The dataset used in this study originates from a community-based CKD screening conducted in the Mirzapur sub-district of Tangail, Bangladesh, a rural and peri-urban region covered by the Mirzapur demographic surveillance system (DSS). Adults aged $\geq 18$ years with at least five years of residency were selected using age-stratified random sampling, yielding 928 enrolled participants, of whom 872 completed all study procedures. Following home-based enrolment, trained community health workers obtained written informed consent, performed interviews and physical examinations, and referred participants to a hospital laboratory for the first assessment, during which blood and urine samples were collected to measure serum creatinine, fasting glucose, lipid profile, hemoglobin, and the urine albumin-to-creatinine ratio (ACR). Participants exhibiting an estimated glomerular filtration rate (eGFR) <60<60 mL/min/1.73 m2 and/or ACR $\geq 30$ mg/g were invited for a second assessment after a 3-month interval to confirm persistent abnormalities. CKD diagnosis was made using the CKD-EPI equation to estimate eGFR and stage following the National Kidney Foundation Kidney Disease Outcomes Quality Initiative (NKF/KDOQI) guidelines [inker2014kdoqi], requiring reduced eGFR and/or albuminuria sustained for at least three months. The study received ethics approval from the ICDDR,B Research Review Committee and Ethical Review Committee (registration no. 19081), and all participants provided written informed consent. Further details on the data collection process are provided in [sarker2021community].
In this study, we focused exclusively on participants with CKD stages 1 to 3, as these stages represent the early phase of kidney impairment when symptoms are typically absent and community-based screening can still play a critical preventive role. Participants in stages 4 and 5 were excluded because these advanced stages are comparatively rare in the dataset and are usually accompanied by clinically evident symptoms, making them less relevant to early screening. After removing records with missing values, the final analytic dataset comprised 284 participants, including 112 CKD (stage 1: n = 25; stage 2: n = 51; stage 3: n = 36), and 172 non-CKD controls. Table 1 presents the descriptive characteristics of the 24 study variables, all of which were transformed into categorical variables for analysis. These variables span multiple groups, including socio-demographic (SD) factors (age group, gender, literacy status, occupation, and marital status), lifestyle and habit (LH) variables (sleep duration, tobacco smoking, and smokeless tobacco use), medical history (MH) (hypertension, diabetes, heart disease, stroke, and family histories of major conditions), and physical or clinical examination (CE) (body mass index (BMI) category, abdominal obesity, undernutrition status, and anemia), and pathology tests (Path) (presence of urinary red blood cells, low serum albumin, hypercholesterolemia, low HDL-cholesterol, and hypertriglyceridemia). These variables collectively represent the multidimensional risk profile relevant to early-stage CKD screening.
Table 1: Primary dataset description
Group Feature Description Value CKD n=112 (%) Non-CKD n=172 (%)
Socio-demographic (SD) Age Age in years 18-30y 3 (2.7) 52 (30.2)
31-39y 12 (10.7) 22 (12.8)
40-49y 15 (13.4) 47 (27.3)
50-60y 33 (29.5) 51 (29.6)
60+y 49 (43.7) 0 (0)
Gender Male or Female Female 73 (65.2) 95 (55.2)
Illiterate Without formal education Yes 49 (43.7) 55 (32)
Occupation Occupation of participant Farmer 8 (7.1) 12 (7)
Housewife 67 (59.8) 80 (46.5)
Others 37 (33) 80 (46.5)
Marital status Participant marital status Married 81 (72.3) 142 (82.6)
Widowed 28 (25) 8 (4.6)
Others 3 (2.7) 22 (12.8)
Lifestyle & habit (LH) Sleeping duration <<7 hours daily sleep (Yes) Yes 46 (41.1) 35 (20.3)
Tobacco smoker Currently smoke any form of tobacco Yes 13 (11.6) 37 (21.5)
Smokeless tobacco Currently consumed any form of smokeless tobacco Yes 40 (35.7) 35 (20.3)
Medical history (MH) Hypertension History of or screening-detected hypertension Yes 75 (67) 20 (11.6)
Diabetes History of or screening-detected diabetes Yes 30 (26.8) 24 (13.9)
Heart disease History of heart disease Yes 11 (9.8) 6 (3.5)
Stroke History of stroke Yes 10 (8.9) 5 (2.9)
Family diabetes Family history of diabetes Yes 28 (25) 37 (21.5)
Family hypertension Family history of hypertension Yes 43 (38.4) 61 (35.5)
Family CKD Family history of CKD Yes 12 (10.7) 14 (8.1)
Clinical examination (CE) BMI Body mass index, underweight (<<18.5), Underweight 10 (8.9) 23 (13.4)
[…]
         obese ($\geq$30) Obese 44 (39.3) 43 (25)
[…]
Undernutrition Mid-upper arm circumference, men<<25cm, women<<24cm Yes 20 (17.9) 36 (20.9)
Anemia Blood hemoglobin, men<<13g/dL, women<<12g/dL Yes 41 (36.6) 33 (19.2)
Pathology tests (Path) RBC Presence of red blood cells in urine Yes 20 (17.9) 0 (0)
Low serum albumin Serum albumin<<3.5 g/dL Yes 12 (10.7) 4 (2.3)
Hypercholesterolaemia Serum cholesterol>>200 mg/dL Yes 28 (25) 27 (15.7)
Low HDL-cholesterol Serum HDL-cholesterol<<40mg/dL Yes 32 (28.6) 63 (36.6)
Hypertriglyceridemia Serum triglyceride>>150mg/dL Yes 48 (42.8) 56 (32.6)
For external validation, we incorporated three publicly available datasets (summarized in Table 1). The UCI-2015 CKD dataset [chronic_kidney_disease_336], collected from a hospital in India, is the most widely used benchmark dataset for machine learning research on CKD detection. It contains clinical records from 400 patients admitted over a two-month period, comprising 250 CKD cases and 150 non-CKD controls. The dataset includes 24 features spanning multiple groups: one socio-demographic variable, one lifestyle and habit-related variable, four medical history variables, four clinical examination measures, and fourteen pathology test results, including two standard pathology tests for CKD diagnosis, such as urinary albumin and serum creatinine. Notably, CKD stages are not reported in this dataset.
Another dataset used for external validation is the UCI-2023 CKD dataset [risk_factor_prediction_of_chronic_kidney_disease_857], collected at Enam Medical College in Savar, Bangladesh. It contains records from 200 patients, including 128 confirmed CKD cases and 72 non-CKD controls, with CKD stages 1–5 explicitly annotated. The dataset comprises 27 features spanning several groups: one socio-demographic attribute, one lifestyle and habit-related variable, four medical history indicators, six clinical examination measurements, and fifteen pathology test results, including three standard pathology tests for CKD diagnosis, such as urinary albumin, serum creatinine, and eGFR.
The third dataset is the TH CKD dataset [Al-Shamsi2018], derived from hospital records at Tawam Hospital in Al-Ain City, Abu Dhabi, and is one of the few Middle Eastern clinical datasets used for CKD risk modeling. It includes 491 adult patients, of whom 56 were identified with moderate-to-advanced CKD (stages 3–5). The remaining 435 participants were labeled as the “non-CKD” category. However, because early-stage CKD (stages 1–2) was not distinguished from healthy controls, this negative class does not constitute a clinically reliable non-CKD population. Consequently, these samples were excluded from cross-validation in our study, and only the confirmed CKD cases were used for external evaluation. The dataset contains 21 features spanning multiple groups: two socio-demographic variables, one lifestyle and habit-related factor, ten medical history attributes, three clinical examination measures, and five pathology test indicators, including two standard pathology tests for CKD diagnosis, such as serum creatinine and eGFR.
2.2 Data Preprocessing
Among the four datasets incorporated into this study, the UCI-2015 CKD dataset [chronic_kidney_disease_336] contained missing values requiring systematic preprocessing. Missing numeric features were addressed using the multivariate imputation by chained equations (MICE) algorithm, a widely adopted multiple-imputation framework that iteratively models each variable with missing data as a function of the other variables in the dataset [vanBuuren2011MICE]. Rather than replacing missing entries with a single deterministic value, MICE generates multiple plausible imputations, thereby preserving the underlying variability and reducing bias associated with single-imputation techniques. For categorical variables, missing entries were imputed using the most-frequent strategy, an approach demonstrated to be effective for maintaining distributional integrity in classification-focused machine learning tasks [adhikari2022comprehensive].
To ensure consistency, shared features across external and primary datasets were identified, harmonized, and renamed. Continuous variables were discretized into categorical, where appropriate, following our dataset’s strategy (Table 1). This normalization aligned heterogeneous sources, enabling model training on the primary dataset and evaluation on external datasets.
Finally, all four datasets were transformed using one-hot encoding to convert nominal categorical variables into a numerical format compatible with machine learning algorithms [pedregosa2011scikit]. This approach avoids spurious ordinal assumptions, improves interpretability, and ensures consistency across the diverse classification models evaluated in this study.
2.3 Feature Ranking and Selection
In this study, all variables were grouped into five conceptually coherent feature groups: socio-demographic (SD), lifestyle and habit (LH), medical history (MH), clinical examination (CE), and pathological tests (Path). These groups were defined based on clinical relevance and their potential utility for early CKD detection. The grouping also reflects their relative availability in community or low-resource settings. For example, pathology test features are typically the least accessible and least affordable in rural or resource-constrained environments, whereas CE features can be obtained by trained personnel using inexpensive portable devices. In contrast, SD, LH, and MH features can be collected through structured interviews without specialized equipment. To further examine the combined predictive utility of these groups, we also evaluated three merged feature groups: SD-LH, SD-LH-CE, and SD-LH-MH-CE.
Beyond group-wise evaluation, we employed a diverse set of feature selection (FS) techniques to identify the most informative subset of predictors across all features (denoted as set S1) and across all features except the pathology tests group (denoted as set S2). Ten complementary FS methods were applied, including a statistical filter technique such as Mann–Whitney U (MWU) test [mann1947test], a regularization-based method such as the least absolute shrinkage and selection operator (LASSO) [tibshirani1996lasso], and wrapper-based recursive approaches such as recursive feature elimination with cross-validation (RFECV) [guyon2002rfe] using a variety of ML estimators, including logistic regression (LR) [hosmer2013applied], decision tree (DT) [breiman1984cart], random forest (RF) [breiman2001rf], gradient boosting (GB) [friedman2001gbm], adaptive boosting (AB) [freund1997adaboost], extra trees (ET) [geurts2006extratrees], extreme gradient boosting (XGB) [chen2016xgboost], and CatBoost (CB) [dorogush2018catboost].
Following the application of these ten feature selection methods, we derived two consolidated feature subsets: (i) an intersection set, representing features consistently selected across all ten FS techniques, and (ii) a union set, comprising the complete set of unique features identified by any of the FS methods. These aggregated sets allow robust comparison of model performance under conservative (intersection) and inclusive (union) feature selection paradigms.
2.4 Machine Learning Classifiers
To develop a robust and generalizable model for early CKD prediction, we employed twelve state-of-the-art ML classifiers spanning diverse algorithmic paradigms. These classifiers were selected based on their demonstrated effectiveness in handling structured, tabular, and high-dimensional clinical data, as well as their widespread use in prior CKD detection research [sanmarchi2023predict, delrue2024application, gogoi2025machine]. The models encompass linear, nonlinear, ensemble-based, and neural network–based learning strategies, thereby ensuring comprehensive exploration of the feature space.
The set includes a linear model such as logistic regression (LR) [hosmer2013applied], a tree-based classifier such as decision tree (DT) [breiman1984cart], a distance-based model such as $k$-nearest neighbours (KNN) [cover1967knn], and a kernel-based algorithm such as support vector machine (SVM) [cortes1995svm]. To capture variance via bootstrap aggregation (bagging), we incorporated random forest (RF) [breiman2001rf] and extra trees (ET) [geurts2006extratrees]. For sequential error-correcting boosting strategies, we employed adaptive boosting (AB) [freund1997adaboost], gradient boosting (GB) [friedman2001gbm], extreme gradient boosting (XGB) [chen2016xgboost], light gradient boosting (LGB) [ke2017lightgbm], and CatBoost (CB) [dorogush2018catboost]. Finally, we included a neural network–based classifier, the multi-layer perceptron (MLP) [rumelhart1986mlp], to evaluate nonlinear representations learned through back-propagation. Together, these twelve classifiers provide a comprehensive benchmark for assessing CKD predictability across various modeling assumptions and complexity levels.
2.5 Model Training
All machine learning models were trained following a rigorous and reproducible experimental pipeline. To ensure robust estimation of generalization performance, we employed stratified 10-fold cross-validation, a widely used model evaluation strategy that mitigates overfitting by repeatedly training and testing models on different partitions of the data while preserving class distribution [kohavi1995cv]. For reproducibility, all random processes (data shuffling, model initialization, and parameter sampling) were controlled using a fixed random_state = 42.
Model development and evaluation were implemented using the scikit-learn library [pedregosa2011scikit], along with specialized optimization frameworks for advanced gradient boosting models. To identify the optimal hyperparameters for each classifier, we utilized Optuna, an efficient, modern hyperparameter optimization framework that employs sequential model-based optimization (SMBO) with adaptive sampling strategies to search high-dimensional parameter spaces [akiba2019optuna]. During optimization, each candidate configuration was evaluated within the cross-validation workflow to ensure unbiased performance estimation. This systematic training and optimization procedure enabled the selection of high-performing, well-calibrated models across all feature subsets and ensured fair comparison among the twelve machine learning classifiers.
2.6 Performance Metrics
Given that the dataset used in this study is imbalanced, with the CKD class representing the minority and clinically significant positive class, we employed evaluation metrics that are robust to class imbalance and that reflect clinically meaningful performance. For binary classification (CKD as the positive class), we report balanced accuracy, sensitivity (recall for CKD), area under the receiver operating characteristic curve (AUC-ROC), F1 score (for CKD), and macro-precision. The definition of those metrics is provided in B. For all metrics, we additionally compute the 95% confidence intervals (CI) across the 10-fold cross-validation splits to quantify statistical uncertainty. This provides a robust estimate of model reliability and generalization performance.
2.7 Explainability
Explainability is a critical component in the development and deployment of machine learning models for clinical applications. It enhances transparency, fosters trust among healthcare professionals, and supports clinical validation by elucidating how input features contribute to model predictions. In the context of CKD detection, explainability enables clinicians to understand which factors most strongly influence an individual’s predicted CKD risk, thereby facilitating more informed decision-making and improving patient management.
Among the different explainability methods, SHAP (SHapley Additive exPlanations) [lundberg2017shap] has emerged as one of the most powerful due to its rigorous theoretical grounding in cooperative game theory. SHAP provides both global interpretability by quantifying overall feature importance across the population and local interpretability, allowing an examination of how specific features contribute to the prediction for an individual patient. These dual capabilities make SHAP especially suitable for clinical environments, where understanding both broad model behaviour and case-specific reasoning is essential. SHAP has also been successfully applied in previous studies on CKD prediction [gogoi2025machine].
Motivated by its demonstrated effectiveness and theoretical strengths, this study employs SHAP to analyze the contribution of each selected feature to the model’s predictions. Through this approach, we aim to bridge the gap between predictive performance and clinical interpretability, advancing the development of a transparent, explainable, and trustworthy AI-driven tool for early CKD detection.
3 Results
3.1 Selected Features
Table 2 summarizes the ranking of features obtained using ten complementary feature selection techniques for both S1 (all features) and S2 (all features excluding pathology tests). A small subset of features consistently appears among the highest-ranked predictors across nearly all methods, most notably Hypertension, Age_60+years, and the pathology-related feature RBC. Their consistently low rank values indicate strong and stable predictive importance, underscoring their critical role in identifying individuals at risk of CKD. Likewise, several clinically recognized risk factors—such as Anemia, Diabetes, Daily sleep<<7h, and BMI are frequently assigned favourable ranks across multiple selection strategies, reinforcing their well-established relevance in CKD detection.
Although some degree of variability is observed across the different methodological families, model-driven approaches (e.g., RFECV), regularization-based techniques (e.g., LASSO), and statistical tests (e.g., Mann-Whitney U (MWU) test), the overall consensus on the top-ranked features highlights the robustness and stability of the feature selection process. Collectively, these consistently identified predictors form a strong, clinically meaningful foundation for developing reliable and generalizable CKD prediction models.
In addition to individual rankings, we derived both common and union feature subsets for S1 and S2. The common subset for S2 comprises Hypertension and Age_60+years, while the S1 common subset additionally includes the pathology-based feature RBC. These shared features represent the most universally selected predictors across all methods, providing compact yet highly informative feature sets for downstream model development.
3.2 Performance of ML Models
To assess the discriminative capability of different feature domains for CKD prediction, we systematically evaluated multiple feature sets, including individual feature groups, their combinations, and the feature subsets derived from S1 and S2 using the ten feature selection techniques described in Section 3.1. This comprehensive comparison enables a detailed understanding of how various domains and selected features contribute to model performance.
Table 3 summarizes the performance of all feature sets, reporting results based on the best-performing model for each set. Every feature set was evaluated independently using the twelve machine learning classifiers introduced in Section 2.4. All models were trained and validated using stratified 10-fold cross-validation, and performance was quantified using appropriate evaluation metrics accompanied by 95% confidence intervals to ensure statistical robustness.
The best-performing model for each feature set was identified primarily based on the highest balanced accuracy, reflecting the need to account for class imbalance in CKD prediction. In cases where two or more models achieved identical balanced accuracy, we applied a secondary ranking strategy prioritizing higher sensitivity for the CKD (positive) class, followed by AUC-ROC, and subsequently the CKD-specific F1 score. This hierarchical selection process ensures that the chosen model not only performs well overall but also aligns with clinical priorities by maximizing the correct identification of CKD cases.
Table 2: Feature sets with the ranking of features (where lower rank values indicate higher predictive importance) for S1 (all features) and S2 (all features excluding the pathology test group), derived using ten diverse feature ranking and selection techniques, highlighting consistent high-impact predictors for early CKD detection.
Feature Group RFECV based feature ranking MWU LASSO
LR RF GB DT AB ET XGB CB
S1 S2 S1 S2 S1 S2 S1 S2 S1 S2 S1 S2 S1 S2 S1 S2 S1 S2 S1 S2
Hypertension MH 3 2 1 1 1 1 1 1 3 2 1 1 3 2 1 1 1 1 3 2
Age_60+y SD 1 1 2 2 2 2 2 2 1 1 2 2 1 1 2 2 2 2 1 1
RBC Path 2 3 3 3 2 3 2 6 3 2
Anemia CE 5 4 4 4 3 6 4 5 3 4 4 7 6 10
Diabetes MH 4 3 5 3 4 4 3 7 4 3 3 11 10
Daily sleep <<7h LH 6 5 8 6 7 6 5 5 4
Age_18-30y SD 5 5 5 4 3 4 3
BMI_Obese CE 5 12 11 6
Smokless tobacco LH 9 8 9
Age_40-49y SD 10 9 8
MaritialS_Widowed SD 5 4
Low serum albumin Path 6 4
Abdominal obesity CE 8 7
Gender SD 8 7
Occup_Housewife SD 11 5
Family hypertension MH 9 12
Stroke history MH 5
BMI_Underweight CE 6
BMI_Normal CE 9
Table 3: Performance of best models for various feature sets. Bold values represent the highest performance metric in each group
Feature set Best Model Performance $\pm$ 95% Confidence Interval (CI)
Balanced accuracy Sensitivity (CKD) AUC-ROC F1 (CKD) Precision (macro)
All CB $84.54_{\pm 3.27}$ $79.55_{\pm 5.94}$ $0.900_{\pm 0.05}$ $81.22_{\pm 4.00}$ $85.30_{\pm 2.91}$
Groups and combinations Socio-Demo (SD) DT $73.27_{\pm 2.75}$ $65.23_{\pm 6.90}$ $0.816_{\pm 0.03}$ $67.08_{\pm 3.69}$ $74.35_{\pm 2.88}$
Life-style/Habbit (LH) LGB $62.90_{\pm 6.23}$ $53.64_{\pm 10.19}$ $0.648_{\pm 0.07}$ $54.09_{\pm 8.20}$ $63.18_{\pm 6.00}$
Medical-history (MH) DT $78.87_{\pm 7.09}$ $75.15_{\pm 9.74}$ $0.805_{\pm 0.06}$ $74.37_{\pm 8.40}$ $78.99_{\pm 7.13}$
Clinical exam (CE) LR $58.77_{\pm 6.48}$ $57.12_{\pm 12.50}$ $0.593_{\pm 0.06}$ $51.68_{\pm 8.62}$ $58.59_{\pm 6.37}$
Pathology tests (Path) LR $64.24_{\pm 5.74}$ $52.80_{\pm 10.14}$ $0.702_{\pm 0.07}$ $55.12_{\pm 7.41}$ $66.13_{\pm 6.51}$
SD-LH AB $73.20_{\pm 2.98}$ $49.92_{\pm 6.45}$ $0.820_{\pm 0.04}$ $63.95_{\pm 5.31}$ $\textbf{83.53}_{\pm 3.95}$
SD-LH-MH LR $\textbf{82.93}_{\pm 3.55}$ $\textbf{78.64}_{\pm 6.71}$ $0.881_{\pm 0.04}$ $\textbf{79.16}_{\pm 4.26}$ $83.40_{\pm 3.22}$
SD-LH-MH-CE CB $81.87_{\pm 4.26}$ $74.17_{\pm 6.72}$ $\textbf{0.887}_{\pm 0.06}$ $77.80_{\pm 5.59}$ $83.21_{\pm 3.96}$
Selected across all features (S1) RFECV+LR AB $86.19_{\pm 3.67}$ $84.02_{\pm 8.40}$ $0.89_{\pm 0.04}$ $82.91_{\pm 4.30}$ $86.35_{\pm 2.74}$
RFECV+RF SVM $88.45_{\pm 5.17}$ $76.89_{\pm 10.34}$ $0.859_{\pm 0.08}$ $86.24_{\pm 6.80}$ $\textbf{93.72}_{\pm 2.56}$
RFECV+GB GB $88.75_{\pm 5.50}$ $78.64_{\pm 10.31}$ $0.920_{\pm 0.07}$ $86.56_{\pm 7.17}$ $92.96_{\pm 3.70}$
RFECV+DT AB $86.19_{\pm 3.67}$ $84.02_{\pm 8.40}$ $0.895_{\pm 0.04}$ $82.91_{\pm 4.30}$ $86.35_{\pm 2.74}$
RFECV+AB AB $86.19_{\pm 3.67}$ $84.02_{\pm 8.40}$ $0.895_{\pm 0.04}$ $82.91_{\pm 4.30}$ $86.35_{\pm 2.74}$
RFECV+ET GB $87.99_{\pm 4.67}$ $75.98_{\pm 9.34}$ $\textbf{0.931}_{\pm 0.04}$ $85.76_{\pm 6.32}$ $93.44_{\pm 2.23}$
RFECV+XGB XGB $88.16_{\pm 6.17}$ $80.38_{\pm 10.77}$ $0.915_{\pm 0.07}$ $85.76_{\pm 7.85}$ $91.11_{\pm 5.32}$
RFECV+CB DT $\textbf{90.40}_{\pm 5.07}$ $\textbf{84.85}_{\pm 7.39}$ $0.906_{\pm 0.05}$ $\textbf{88.77}_{\pm 6.43}$ $92.14_{\pm 4.99}$
LASSO LR $84.99_{\pm 3.82}$ $82.20_{\pm 9.09}$ $0.910_{\pm 0.04}$ $81.46_{\pm 4.40}$ $85.56_{\pm 3.17}$
MWUtest GB $85.99_{\pm 4.26}$ $79.47_{\pm 7.49}$ $0.925_{\pm 0.05}$ $83.09_{\pm 5.42}$ $87.91_{\pm 4.38}$
Union CB $85.88_{\pm 3.47}$ $80.45_{\pm 6.47}$ $0.898_{\pm 0.05}$ $82.88_{\pm 4.35}$ $87.00_{\pm 3.24}$
Common AB $86.19_{\pm 3.67}$ $84.02_{\pm 8.40}$ $0.895_{\pm 0.04}$ $82.91_{\pm 4.30}$ $86.35_{\pm 2.74}$
Selected across all except pathology tests (S2) RFECV+LR AB $83.50_{\pm 3.53}$ $78.64_{\pm 7.38}$ $0.861_{\pm 0.04}$ $79.77_{\pm 4.41}$ $84.21_{\pm 2.80}$
RFECV+RF MLP $86.25_{\pm 4.79}$ $77.73_{\pm 10.57}$ $0.914_{\pm 0.05}$ $83.14_{\pm 5.92}$ $89.56_{\pm 3.50}$
RFECV+GB CB $\textbf{89.23}_{\pm 6.08}$ $\textbf{83.11}_{\pm 9.13}$ $0.906_{\pm 0.07}$ $\textbf{87.12}_{\pm 7.60}$ $\textbf{90.92}_{\pm 5.74}$
RFECV+DT DT $87.75_{\pm 7.44}$ $79.55_{\pm 11.52}$ $0.913_{\pm 0.07}$ $85.11_{\pm 9.57}$ $90.19_{\pm 6.77}$
RFECV+AB AB $83.50_{\pm 3.53}$ $78.64_{\pm 7.38}$ $0.861_{\pm 0.04}$ $79.77_{\pm 4.41}$ $84.21_{\pm 2.80}$
RFECV+ET CB $86.25_{\pm 4.79}$ $77.73_{\pm 10.57}$ $0.909_{\pm 0.05}$ $83.14_{\pm 5.92}$ $89.56_{\pm 3.50}$
RFECV+XGB RF $87.17_{\pm 7.05}$ $79.55_{\pm 11.52}$ $\textbf{0.921}_{\pm 0.06}$ $84.24_{\pm 8.99}$ $89.30_{\pm 6.27}$
RFECV+CB MLP $86.25_{\pm 4.79}$ $77.73_{\pm 10.57}$ $0.914_{\pm 0.05}$ $83.14_{\pm 5.92}$ $89.56_{\pm 3.50}$
LASSO LR $83.50_{\pm 3.53}$ $78.64_{\pm 7.38}$ $0.892_{\pm 0.04}$ $79.77_{\pm 4.41}$ $84.21_{\pm 2.80}$
MWUtest CB $84.42_{\pm 6.61}$ $78.64_{\pm 10.07}$ $0.896_{\pm 0.07}$ $80.84_{\pm 7.74}$ $85.61_{\pm 5.63}$
Union LGB $84.58_{\pm 6.60}$ $79.62_{\pm 11.08}$ $0.898_{\pm 0.06}$ $80.95_{\pm 8.19}$ $85.52_{\pm 5.95}$
Common AB $83.50_{\pm 3.53}$ $78.64_{\pm 7.38}$ $0.861_{\pm 0.04}$ $79.77_{\pm 4.41}$ $84.21_{\pm 2.80}$
Overall, the results in Table 3 highlight that models trained on the selected features from the full feature set (S1) achieve the strongest predictive performance. The highest balanced accuracy of 90.40% is obtained using a DT model applied to the RFECV with CatBoost selected S1 feature subset (Table 1), achieving the highest sensitivity of 84.85% and F1 score of 88.77%. Among the individual feature groups, the Medical History (MH) and Socio-Demographic (SD) sets demonstrate competitive predictive ability, with balanced accuracies of 78.87% and 73.27%, respectively. Notably, the combined SD-LH-MH set outperforms all individuals and combined groups, achieving a balanced accuracy of 82.93%, which is higher than when clinical examination features are additionally included (i.e., SD-LH-MH-CE).
Feature subsets derived through feature selection methods (for both S1 and S2) also exhibit strong predictive performance. Across both settings, RFECV-based feature selection consistently outperforms LASSO and the MWU test, highlighting the advantage of model-driven wrapper approaches. Importantly, the selected feature set from S2, which excludes pathology-related variables, achieves performance close to that of the full feature set. Using a CatBoost model on the RFECV+GB-selected S2 feature subset (Table 1), we obtain a balanced accuracy of 89.23%, sensitivity of 83.11%, F1 score of 87.12%, and precision of 90.92%. These values are only 1–2% lower than the best-performing S1 models, demonstrating that CKD can be predicted with competitive accuracy using solely non-pathological, non-laboratory features. Figures 2 and 3 together illustrate the comparative performance and classification behaviour of models trained on those feature sets. This finding underscores the feasibility of deploying CKD screening models in community-based and low-resource settings without the need for laboratory or pathology tests.
BalancedAccuracySensitivity (CKD)AUCROCF1 (CKD)Precision0.60.60.70.70.80.80.90.911Performance scoreAllBest (S2)Best (S1)
Figure 2: Performance comparison of machine learning models trained on three different feature configurations: the full feature set, the best-performing subset excluding pathology features (S2), and the best-performing subset across all feature domains (S1). Error bars represent 95% confidence intervals computed over 10-fold cross validation.
Figure 3: Confusion matrices illustrating the prediction results for CKD and non-CKD cases using three feature configurations: (a) the full feature set, (b) the best-performing subset without pathology features, and (c) the best-performing subset including all features. Values represent both the counts and percentages of true positives, true negatives, false positives, and false negatives.
3.3 Comparison with State-Of-The-Art Tools
Table 4 provides a comprehensive comparison between the proposed CKD prediction models and several state-of-the-art (SOTA) CKD screening tools, including SCORED [bang2007development], kshirsagar2008simple (simplified version), thakkinstian2011simple, kearns2013predicting, and kwon2019simple. For each tool, the table reports balanced accuracy, sensitivity (CKD), AUC-ROC, F1 score (CKD), and macro-precision. The two proposed models, one using the best-performing S1 feature set with a Decision Tree, and the other using the best S2 subset with CatBoost, demonstrate substantially higher balanced accuracy, sensitivity, F1 score, and precision than all existing tools.
Table 4: Performance comparison with state-of-the-art (SOTA) CKD screening tools. The SPS tool does not directly classify individuals as CKD or non-CKD; instead, it assigns four risk categories: low, intermediate-low, intermediate-high, and high. Accordingly, the low category was treated as non-CKD, while the remaining three categories were treated as CKD for evaluation. Asterisks following each performance value indicate the results of pairwise statistical comparisons between the proposed model and the baseline tools.
Tool Balanced accuracy Sensitivity (CKD) AUC-ROC F1 (CKD) Precision (macro)
SCORED [bang2007development] 76.35(∗∗∗S1,∗∗∗S2) 66.07(∗∗∗S1,∗∗∗S2) 0.8507(∗S1,∗∗∗S2) 70.81(∗∗∗S1,∗∗∗S2) 77.98(∗∗S1,∗∗∗S2)
[…]
kwon2019simple 77.49(∗∗∗S1,∗∗∗S2) 64.29(∗∗∗S1,∗∗∗S2) 0.8537(∗S1,∗∗S2) 72.00(∗∗∗S1,∗∗∗S2) 80.71(∗∗S1,∗∗∗S2)
[…]
This study (best S1 with DT) 90.40 84.85 0.9062 88.77 92.14
This study (best S2 with CB) 89.23 83.11 0.9062 87.12 90.92
*$p<0.05$, **$p<0.01$, ***$p<0.001$
Asterisks following each performance value indicate the statistical significance of improvements relative to the SOTA methods, with comparisons presented in a fixed order (S1, S2). Most comparisons show highly significant (**$p<0.01$) or extremely significant (***$p<0.001$) improvements, underscoring that the proposed models offer markedly superior predictive performance over current screening tools. Overall, the table highlights the strong advantage of our feature selection strategy and machine learning framework in achieving more accurate and clinically meaningful CKD screening.
3.4 External Validation
To evaluate the applicability and generalizability of the best-performing S1 and S2 feature subsets, we conducted external validation using three independent publicly available datasets: UCI-2023 [chronic_kidney_disease_336], UCI-2015 [risk_factor_prediction_of_chronic_kidney_disease_857], and TH [Al-Shamsi2018]. For each of these datasets, we first identified the set of features shared with our dataset and harmonized the naming conventions to ensure consistency across sources. Based on these harmonized features, we then constructed the S1 and S2 subsets by selecting the variables that overlapped with our optimized S1 and S2 feature sets, respectively. The resulting common, S1, and S2 feature subsets for each external dataset are summarized in Table 5. Notably, the TH dataset did not contain any pathology-related features that matched those in our dataset, and therefore, its S1 subset could not be constructed. This process enables a direct and fair assessment of how well the models trained on our dataset generalize when applied to diverse populations and measurement contexts.
Table 5: Feature sets for external validation
Dataset Name Features
UCI-2023 [chronic_kidney_disease_336] Common Hypertension, Age_60+y, Diabetes, Anemia, RBC, Age_31-39y
S1 subset Hypertension, Age_60+y, Diabetes, Anemia, RBC
S2 subset Hypertension, Age_60+y, Diabetes, Anemia, Age_31-39y
UCI-2015 [risk_factor_prediction_of_chronic_kidney_disease_857] Common Hypertension, Age_60+y, Diabetes, Anemia, RBC, Age_18-30y,
Age_31-39y, Age_40-49y, Age_50-60y
S1 subset Hypertension, Age_60+y, Diabetes, Anemia, RBC
S2 subset Hypertension, Age_60+y, Diabetes, Anemia, Age_31-39y
TH [Al-Shamsi2018] Common Hypertension, Age_60+y, Diabetes, BMI_Obese, Gender, Age_31-39y,
Age_18-30y, Age_40-49y, Age_50-60y, Heart disease,
Tobacco smoker, Hypercholesterolemia, Hypertriglyceridemia
S1 subset None common pathology tests
S2 subset Hypertension, Age_60+y, Diabetes, Age_31-39y
Table 6 presents the external validation performance of the machine learning models trained on our dataset and evaluated on the three independent datasets. Across all datasets, both the common and subset-based models demonstrate strong sensitivity (e.g., 80.47% in UCI-2023) and competitive balanced accuracy (e.g., 79.33% in UCI-2015), indicating good generalizability of the trained models to external populations. Notably, the TH dataset contains only CKD cases; therefore, only sensitivity is reported for this cohort, with the S2 subset demonstrating exceptional performance at 98.21%.
Table 6: External validation performance of machine learning models trained on our dataset and tested on three external datasets.
Test dataset Feature set Balanced accuracy Sensitivity (CKD) AUC-ROC F1 (CKD) Precision (macro)
UCI-2023 [chronic_kidney_disease_336] Common 76.35 80.47 0.819 82.07 75.64
S1 subset 76.35 80.47 0.844 82.07 75.64
S2 subset 74 75.78 0.804 79.18 72.78
UCI-2015 [risk_factor_prediction_of_chronic_kidney_disease_857] Common 78.93 77.2 0.826 81.78 77.46
S1 subset 79.33 78 0.881 82.28 77.9
S2 subset 77.13 73.6 0.842 79.48 75.55
TH [Al-Shamsi2018] Common 96.43
S2 subset 98.21
Across all datasets, models using the S1 subset (for UCI-2023 and UCI-2015) and the S2 subset (for TH) consistently outperform those using the common feature set, demonstrating improved discriminatory power. This consistent improvement underscores the effectiveness of the feature selection strategy developed in this study and supports its generalizability and applicability across diverse datasets.
3.5 Model Interpretation and Key Predictors
To enhance model transparency and clinical interpretability, we applied SHAP to analyze the contributions of the selected features in the best-performing model. Specifically, SHAP values were computed for the feature subset obtained using the RFECV+CB approach within the S1 setting, and interpreted using our best-performing DT classifier for CKD detection. Figure 4 shows the resulting SHAP summary plot, where features are ranked along the Y-axis according to their mean absolute SHAP values, reflecting their overall importance and average impact on model predictions. The X-axis displays the SHAP values for individual instances, indicating both the magnitude and direction of each feature’s contribution. Positive SHAP values push the prediction toward CKD, whereas negative values push it toward non-CKD. Each point represents a single data instance, with colour denoting the magnitude of the one-hot encoded feature value (red indicating presence/“yes” and blue indicating absence/“no”). This explainability analysis provides insight into how key predictors influence decision-making within the model, supporting clinical trust and interpretability.
Figure 4: SHAP summary plot illustrating the contribution of the best-performing S1 feature set to the Decision Tree model’s predictions for CKD detection.
Figure 5 presents SHAP waterfall plots for two representative prediction instances, a correctly classified CKD (Figure 5(a)) and non-CKD (Figure 5(b) case. Each plot illustrates how individual feature values contribute to shifting the model output from the baseline prediction (the expected value of the model) toward the final predicted probability. In Figure 5(a), several high-risk indicators, including Age_60+y, Hypertension, and Gender, push the prediction strongly toward CKD, resulting in a final output close to 1. Features with negative contributions (e.g., absence of BMI_Obese, Anemia, RBC) slightly offset the prediction but do not reverse the decision. In Figure 5(b), the absence of key CKD risk factors (e.g., Hypertension, Age_60+y, RBC) drives the prediction toward the non-CKD class, producing a low predicted value. Some features, such as Family hypertension contribute small positive shifts, but the strong negative contributions dominate, keeping the final prediction below the decision threshold.
Figure 5: SHAP waterfall plot for a correctly classified (a) CKD and (b) Non-CKD case, illustrating the feature values contributing to the model prediction.
4 Discussion
4.1 Insights and Observations
This study provides several methodological and clinical insights that strengthen the evidence for using minimal, non-invasive features to develop robust and generalizable models for early CKD detection.
First, our findings show that the RFECV-based feature selection method consistently identifies smaller yet highly discriminative feature subsets compared with LASSO and the Mann–Whitney U test. This observation is important because it demonstrates that wrapper-based approaches such as RFECV, particularly when combined with strong base learners such as CatBoost, are more effective in isolating the most informative predictors without inflating the feature space. The smaller feature subsets produced by RFECV preserve predictive strength while reducing redundancy, which is critical for practical deployment in community screening settings.
Second, the results reaffirm that adding more features does not necessarily improve classification performance. For S1, the common feature subset (Hypertension, Age_60+y, and RBC) outperforms both the union and all-feature sets, despite the latter containing considerably more variables. Similarly, for S2, the common subset (Hypertension, Age_60+y) surpasses the performance of the largest S2 feature set (SD–LH–MH–CE). These findings emphasize that a compact set of well-chosen, clinically relevant features can outperform larger, noisier combinations. This has direct implications for model interpretability, ease of data collection, and feasibility of large-scale screening in low-resource environments.
Third, the common feature subsets themselves deliver remarkably strong performance. In S1, the three-feature subset (Hypertension, Age_60+y, RBC) achieves a balanced accuracy of 86.19%, which is highly competitive and sufficient for clinical triage despite its minimal size. Even more striking is the S2 common subset (Hypertension and Age_60+y), which achieves 83.50% balanced accuracy while requiring only non-pathological, easily obtainable features. This highlights the significant potential for deploying CKD risk assessment tools in settings lacking laboratory infrastructure, where pathology tests may be inaccessible.
Fourth, while no single classifier consistently dominates across all feature sets, we observe that many models achieve closely comparable performance, reflecting the stable predictive structure in the data. However, AdaBoost emerges as the most frequently top-performing classifier (Figure 1). This suggests that boosting-based ensemble methods may be particularly well-suited for integrating heterogeneous, mixed-type clinical features when predicting CKD.
Fifth, although our focus was on early-stage CKD detection (stages 1–3), the model demonstrates excellent performance in detecting advanced-stage CKD as well. The TH dataset, which comprises only CKD stages 3–5, achieves an exceptional sensitivity of 98.21% using the S2 feature subset. This robustness across disease stages suggests that the selected features capture fundamental CKD risk patterns regardless of severity. Moreover, successful external validation on datasets from India (UCI-2015) and the UAE (TH) indicates that our approach generalizes reasonably well across ethnically diverse populations in South Asia and the Middle East.
Finally, the comparison with state-of-the-art CKD screening tools reveals that our models not only outperform existing methods in accuracy, sensitivity, and AUC-ROC but also rely on fewer and more readily available features. Tools such as SCORED, kshirsagar2008simple, kshirsagar2008simple, and kearns2013predicting often require broader or more specialized input variables (see Table 2), some of which may not be accessible in low-resource settings. In contrast, our optimized S1 and S2 subsets achieve superior performance with simpler, more scalable feature requirements, highlighting the practical advantage of our methodology for population-level CKD screening.
4.2 Limitations and Future Work
Although this study demonstrates strong predictive performance and compelling generalizability for early CKD detection, several limitations should be acknowledged.
First, the primary dataset used for model development was collected from a specific geographic region and demographic profile, which may not fully represent broader global populations. While external validation using datasets from India and the UAE supports cross-population generalizability, further evaluation on datasets from Africa, Europe, and Latin America would strengthen evidence of worldwide applicability.
Second, the dataset size was relatively small; however, this is the first CKD screening study conducted following the standard CKD diagnosis technique in Bangladesh and, more broadly, across South Asia. Although large datasets such as the UK Biobank or the U.S. CRIC cohort exist, they represent substantially different ethnic groups and are therefore unlikely to reflect South Asian patient characteristics. Moreover, such datasets are not readily accessible. The three external datasets we used also lacked complete alignment with our feature space and CKD staging structure. For example, none of them included the full set of best-performing features identified in our study, limiting the ability to perform comprehensive cross-validation. Additionally, the TH dataset contained only CKD stage 3–5 cases and no non-CKD participants, restricting the calculation of performance metrics that require both classes (e.g., balanced accuracy, AUC). These constraints highlight the urgent need for more complete, high-quality CKD datasets, particularly in underrepresented regions, to support broader validation.
Third, the cross-sectional nature of the datasets prevents evaluation of model performance in predicting CKD progression or transitions between disease stages. Longitudinal datasets are essential for developing models that not only detect CKD early but also forecast disease trajectories and inform personalized intervention strategies.
Fourth, although the primary dataset was imbalanced, we elected not to apply oversampling techniques, as these would introduce synthetic data that may not reflect true clinical patterns. Instead, we employed the “balanced” class-weight parameters available in several machine learning algorithms. While this preserves data authenticity, future studies may explore advanced imbalance-handling methods that avoid synthetic data generation, such as focal loss or cost-sensitive learning.
Fifth, despite evaluating a comprehensive set of machine learning classifiers, our analysis focused on classical ML methods. Although deep learning does not always outperform classical approaches for tabular clinical data, future work could examine hybrid neural models, attention-based architectures, or representation-learning strategies capable of capturing more complex feature interactions.
Finally, this study has not yet undergone clinical deployment. Prospective validation in real-world clinical or community settings will be a crucial next step to evaluate usability, acceptability, workflow integration, and impact on early CKD detection outcomes. Such studies will also help identify behavioural, operational, or socio-cultural factors that influence adoption.
5 Conclusion
This study presents a comprehensive and data-driven framework for early CKD detection using machine learning, demonstrating that high predictive performance can be achieved with minimal, non-invasive, and readily obtainable features. Through rigorous feature selection, extensive model evaluation, comparison with state-of-the-art tools, and validation across three independent external datasets, our findings highlight the robustness, generalizability, and clinical relevance of the proposed approach. Notably, the S2 feature subset, requiring only two non-laboratory variables, achieves performance comparable to full feature sets, underscoring its potential for use in community-based and resource-limited settings where laboratory testing is not feasible. The explainability analysis further enhances transparency and trust by revealing clinically meaningful feature contributions. Overall, this work provides a strong foundation for developing scalable CKD screening solutions and paves the way for future advancements, including longitudinal modeling, integration with digital health platforms, and prospective evaluation in real-world clinical environments.
Summary Table
What was already known on the topic:
  1. 1.
    Early-stage chronic kidney disease (CKD) is typically asymptomatic, and conventional diagnosis relies on pathology tests such as serum creatinine, eGFR, and ACR—tests that are often inaccessible or unaffordable in many low-resource settings.
  2. 2.
    Several community-based CKD screening tools exist (e.g., SCORED), but most were developed in high-income or non–South Asian populations, or primarily target CKD stage $\geq 3$ rather than early-stage disease.
  3. 3.
    Existing machine learning studies on CKD have relied heavily on the UCI-2015 dataset and/or commonly include pathology-test-based features, including serum creatinine or eGFR, introducing information leakage or making them unsuitable for community-level screening in resource-limited environments.
What this study adds to our knowledge:
  1. 1.
    Developed the first community-based, early-stage CKD screening framework tailored to Bangladesh and the broader South Asian context, using only non-laboratory features that are feasible to collect in rural and peri-urban settings.
  2. 2.
    Established a rigorous feature selection pipeline using ten complementary methods, leading to robust, generalizable, and easily obtainable features that can achieve strong predictive performance, outperforming larger feature sets and several existing clinical screening tools, including World Health Organization (WHO)-endorsed SCORED.
  3. 3.
    Conducted extensive external validation across three independent datasets (from India, UAE, Bangladesh), showing the strong generalizability of the selected feature sets and demonstrating applicability across diverse ethnic and geographic populations.
  4. 4.
    Provided transparent model interpretability using SHAP, revealing clinically meaningful feature contributions and increasing trustworthiness for potential deployment in community screening programmes.
CRediT authorship contribution statement
Muhammad Ashad Kabir: Conceptualization, Methodology, Software, Investigation, Validation, Formal analysis, Visualization, Project administration, Writing - Original Draft, Writing - Review & Editing. Sirajam Munira: Investigation, Formal analysis, Validation, Writing - Original Draft. Dewan Tasnia Azad: Conceptualization, Investigation, Validation, Writing - Review & Editing. Saleh Mohammed Ikram: Conceptualization, Investigation, Validation, Writing - Review & Editing. Mohammad Habibur Rahman Sarker: Conceptualization, Data Curation, Validation, Writing - Review & Editing. Syed Manzoor Ahmed Hanifi: Conceptualization, Project administration.
Declaration of competing interest
The authors declare that they have no known competing interests that could have appeared to influence the work reported in this paper.
Data Availability Statement
Of the four datasets used in this study, three are publicly available and can be accessed through their respective repositories as cited in the manuscript. The ICDDR,B dataset cannot be made publicly available due to ethical and institutional restrictions.
Appendix A Dataset summary
Table 1: Summary of CKD datasets used in this study
Dataset Country Sample size CKD Stage Features
CKD Non-CKD Total SD LH MH CE Path Path* Total
UCI-2015 [chronic_kidney_disease_336] India 250 150 400 1 1 3 5 12 2 24
UCI-2023 [risk_factor_prediction_of_chronic_kidney_disease_857] Bangladesh 128 72 200 1–5 1 1 4 6 12 3 27
TH [Al-Shamsi2018] UAE 56 56 3–5 2 1 10 3 3 2 21
This study Bangladesh 112 172 284 1–3 5 3 7 4 5 5 29
*Gold-standard pathology tests for CKD diagnosis, such as serum creatinine, urinary albumin, urinary creatinine, eGFR, and uACR
Appendix B Performance metrics definition
Sensitivity (Recall)
Sensitivity measures the proportion of actual CKD cases correctly identified by the classifier. It is a critical metric in medical diagnostics because missing positive cases can lead to delayed or missed treatment. Sensitivity is defined as:
$\text{Sensitivity}_{\text{CKD}}=\frac{TP}{TP+FN},$ (1)
where $TP$ and $FN$ denote true positives and false negatives, respectively.
Balanced Accuracy
Accuracy alone is misleading for imbalanced datasets. Balanced accuracy mitigates this issue by averaging the recall of both positive and negative classes, ensuring performance assessment independent of class distribution. It is defined as:
$\text{Balanced Accuracy}=\frac{1}{2}\left(\frac{TP}{TP+FN}+\frac{TN}{TN+FP}\right),$ (2)
where $TN$ and $FP$ denote true negatives and false positives, respectively.
F1 Score (CKD)
The F1 score emphasises the harmonic balance between precision and recall for the CKD class, making it suitable when false negatives and false positives are both undesirable. It is computed as:
$\text{F1}_{\text{CKD}}=2\times\frac{\text{Precision}_{\text{CKD}}\times\text{Sensitivity}_{\text{CKD}}}{\text{Precision}_{\text{CKD}}+\text{Sensitivity}_{\text{CKD}}},$ (3)
where
$\text{Precision}_{\text{CKD}}=\frac{TP}{TP+FP}.$ (4)
Precision (macro)
Macro precision computes the unweighted mean precision across both classes, giving equal importance to CKD and non-CKD regardless of class size. This provides an unbiased measure of per-class correctness in the presence of imbalance:
$\text{Precision (macro)}=\frac{1}{2}\left(\text{Precision}_{\text{CKD}}+\text{Precision}_{\text{Non-CKD}}\right).$ (5)
where
$\text{Precision}_{\text{Non-CKD}}=\frac{TN}{TN+FN}.$ (6)
AUC-ROC
The area under the receiver operating characteristic curve evaluates the classifier’s ability to discriminate between CKD and non-CKD cases across all possible decision thresholds. AUC-ROC is threshold-independent and robust to class imbalance, making it a widely recommended metric in clinical risk prediction. Formally,
$\text{AUC-ROC}=\int_{0}^{1}TPR(FPR^{-1}(x))\,dx,$ (7)
where $TPR$ is the true positive rate and $FPR$ is the false positive rate.
Appendix C Extended Results
Table 1: Best-performing feature sets identified for S1 (all features) and S2 (all features excluding pathology tests), based on the highest balanced accuracy achieved across all models.
Best set Features Description Selection approach Best model
S1 Hypertension, Age_60+y, Diabetes, Anemia, BMI_Obese, RBC, Daily sleep <<7h, Gender, Family hypertension Best performing feature set among all features including pathology tests RFECV+CB DT
S2 Hypertension, Age_60+y, Anemia, Diabetes, Daily sleep <<7h, Age_18-30y Best performing feature set among all features excluding pathology tests RFECV+GB CB
01122334455667788Adaptive Boosting (AB)CatBoost (CB)Logistic Regression (LR)Decision Tree (DT)Gradient Boosting (GB)Light Gradient Boosting (LGB)Multilayer Perceptron (MLP)Support Vector Machine (SVM)eXtreme Gradient Boosting (XGB)Random Forest (RF)88665544332222111111CountMachine learning classifier
Figure 1: Distribution of best-performing machine learning models identified across all experiments using different feature subsets. The counts reflect how often each classifier achieved the highest balanced accuracy within a given feature configuration.
Table 2: Features comparison with state-of-the-art (SOTA) CKD screening tools
Tool Country CKD stage Features
Socio-demo Lifestyle & habit Medical history Clinical exam Pathology
SCORED [bang2007development] USA 3–5 Age, Gen HT, DM, PVD, CVD, HF Ane, Ptn
kshirsagar2008simple USA 3–5 Age, Gen HT, DM, PVD, CVD, HF Ane
thakkinstian2011simple Thailand 1–5 Age DM, KS Ane
kearns2013predicting England 3–5 Age, Gen, Eth, Dep Smk HT, DM, PVD, CVD, HF, IHD, Str BP
kwon2019simple S. Korea 3–5 Age, Gen HT, DM, CVD Ane, Ptn
This study (S1) Bangladesh 1–3 Age, Gen Daily sleep HT, DM, Family HT Ane, BMI RBC
This study (S2) Bangladesh 1–3 Age Daily sleep HT, DM Ane
Gen - Gender, Eth - Ethnicity, Dep - Deprivation Score, PVD - Peripheral Vascular Disease, CVD - Cardiovascular Disease, Ane - Anemia, BP - Blood Pressure, Ptn - Dipstick proteinuria, HF - Heart Failure, KS - Kidney Stone, IHD - Ischaemic Heart Disease, Str - Stroke, HT - Hypertension, DM - Diabetes Mellitus, Smk - Smoking, BMI - Body Mass Index, RBC - Presence of red blood cells in urine, Daily sleep - Daily sleep <7<7 hours